# Quantum-private distributed sensing

### v1

@@ I Introduction
Quantum networks allow connected nodes to perform tasks such as multi-user quantum cryptography [1, 2], distributed quantum computation [3, 4], and distributed quantum sensing [5, 6, 7, 8]. The increased connectivity and sharing of resources over networks will likely prompt privacy-enabled protocols to protect sensitive information. A well known example of a private computing protocol is blind quantum computing, which allows clients to access remote quantum processors without revealing the computational algorithm to the servers [9, 3, 4]. Another established application is anonymous quantum communication, where users share cryptographic keys within a quantum network without revealing the identities of the keyholders to an eavesdropper, or other users within the network [10, 11, 12, 13, 14, 15]. Similar ideas from quantum cryptography can be incorporated into remote sensing tasks operating over quantum networks to provide both security and privacy features [16, 17, 18, 19].
Distributed quantum sensing typically involves spatially separated quantum sensors to monitor a global property, i.e., non-localised phenomena, such as magnetic fields, temperature sensing, or drifts between remote clocks [6]. The goal is to estimate a global value $\phi$ which is a linear function of local sensor values $\theta_{i}$, such as the weighted sum—this differs from multi-parameter estimation where each $\theta_{i}$ is estimated, for which it is often optimal to use local strategies [20, 21]. In certain scenarios, distributed sensing can exploit entanglement generation in networks to estimate the global parameter directly with Heisenberg-limited precision scaling [16, 22, 8]. Additionally, entanglement can be used to guarantee security, ensuring only authorised users learn the estimated parameters, and privacy—delegating the sensing task to an untrusted node to perform measurements without learning the parameters [17, 18]. Recently, a two-user scheme has been proposed [23] and demonstrated [24] by sharing a Bell pair between two parties, one who performs the sensing task while only the other learns the estimated parameter.
@@
Here we consider multi-user quantum-private distributed sensing, a task known as private parameter estimation (PPE), for which security bounds are recently established in Ref [25]. We illustrate the conceptual arrangement of the PPE task in Fig.1. The $n$ remote sensors are used to estimate a global phase, $\phi$, which is a sum $\phi\doteq\sum_{i}\theta_{i}$ of $\theta_{i}$ local phase values, with $i\in\{1,\dots,n\}$. The goal is to prevent information about the local $\theta_{i}$ from being learned by any node or eavesdroppers monitoring the communication. Analogous to quantum conference key agreement [26, 27, 15, 28], the correlations of the $n$-partite GHZ state, |$\ket{GHZ}\doteq(\ket{0}^{\otimes{n}}+\ket{1}^{\otimes{n}})/\sqrt{2}$ where |$\ket{0}$ and |$\ket{1}$ are computational basis states, are exploited to perform the distributed sensing task privately. The nodes must first verify that they are sharing the GHZ state to ensure they are performing the protocol in private and with integrity. Once this is achieved, the $\theta_{i}$ are encoded onto the shared GHZ state by coupling each node’s qubit to their respective quantum sensor, then each node measures their qubit and announces the result.
We now outline the state verification protocol [25]. One of the $n$ nodes is designated as the verifier, who is assumed to be honest and has pairwise private channels with each remaining node. An untrusted quantum server prepares and distributes $N_{t}$ copies of the $n$-partite GHZ state to all nodes who store them locally in quantum memories. The verifier randomly chooses half of the copies ($N_{t}/2$) and instructs the group to measure them according to one of the stabilizers, $K_{i}$. The set of stabilizers are given by cyclic permutations of $K_{i}=-X^{(1)}X^{(2)}\dots Y^{(i)}Y^{(i+1)}\dots X^{(n-1)}X^{(n)}$ for $i=1,2,\dots,n$ and $K_{i=n+1}=X^{\otimes n}$. By definition, the measurement outcomes of $K_{i}$ result in a $+1$ eigenvalue for the $n$-qubit GHZ state. Each node performs their respective measurement then communicates their outcome to the verifier directly using their private channels. The verifier computes a failure rate, $f$, as the fraction of copies corresponding to a $-1$ eigenvalue. Provided $f$ is small, one of the remaining untested copies is chosen at random as the target copy, $\rho_{j}$, and is used for the parameter estimation. This verification scheme is closely related to the one in Ref. [29], which is generalised to verify graph states using the full set of stabilizer measurements requiring $2n$ measurements, while in Ref. [25] only $n+1$ measurements are required for the GHZ state. This verification protocol guarantees a lower bound on the fidelity $F(\rho_{j},\ket{GHZ})$ of the target copy with the GHZ state up to some nominal probability,
@@
$\mathbb{P}(\textbf{A})$ is the probability that A is true, given the failure rate $f$, number of users $n$, and two positive variables $c$ and $m$ which relate to the statistical certainty based on the total number of copies, $N_{t}=\lceil 2mn^{5}\text{log}(n)\rceil$, used in the protocol (more details in the Appendix 2). Minimising $c$ improves the fidelity lower bound but decreases the probability of the inequality holding true. The parameter $m$ can be increased commensurately to offset the reduction by $c$, but this requires more test copies. Use of quantum memories in the verification step safeguards against an adversary swapping out non-test copies with non-private states, $|+\rangle^{\otimes{n}}$ where $|+\rangle\doteq(|0\rangle+|1\rangle)/\sqrt{2}$ which allows the adversary to learn the local phases from the outcomes. We assume the future availability of such quantum memories and prepare multi-partite GHZ states in the lab to investigate the implementation of the verification scheme presented in Ref. [25].
@@ II Results
To implement the PPE protocol, the three nodes perform the verification protocol by measuring the stabilizers $K_{i}$, which we pick from a subset of the recorded QST measurements. For the optimal state we obtain an average failure rate $f=0.039\pm 0.005$ and from Ref [25] the verification step is successful if $f\leq 1/(2n^{2})$, in our three-node case this sets a threshold of $f\leq 0.05\dot{5}$ which our state satisfies. Using $f$ we establish a lower-bound fidelity of $0.769\pm 0.006$, via $F(\rho,GHZ)\geq 1-2nf$ when assuming infinite resource-state copies ($c\rightarrow 0$), which is smaller than the fidelity obtained via QST. For our three-node scenario it would be more optimal to use QST, however the measurements scale exponentially with $n$, while the verification protocol scales linearly requiring $n+1$ stabilizers. We also establish an upper-bounded privacy parameter, $\varepsilon_{p}\leq 1.3\pm 0.2$ using $\varepsilon_{p}\leq 24n^{-2}\sqrt{2nf}$, again assuming infinite copies, see Appendix 2 for details. For PPE, $\varepsilon_{p}=0$ corresponds to perfect privacy, wherein $\theta_{i}$ cannot be learned by any node nor an eavesdropper and $\phi$ can be estimated ideally [25]. Using the density matrix reconstructed from QST, we calculate the QFI and obtain $\varepsilon_{p}=0.005\pm 0.002$, which is much smaller than the upper bound from the verification protocol—we will follow up on this in the next section. We implement the phase sensing task for two global phases $\{\phi_{1},\phi_{2}\}$, see Fig. 3. To ensure each phase element encodes $\theta_{i}$ correctly we perform a full phase sweep for each, see Appendix 2 for details, then we independently set $\theta_{i}$ to prepare each $\phi$. To estimate $\phi$, all nodes measure in $X$ then compute the expectation value using all outcomes and evaluate $\phi=\textrm{cos}^{-1}[\langle X^{\otimes n}\rangle]$. The accuracy of the estimate $\hat{\phi}$ to the true value $\phi$ is evaluated as the distance, $\|\hat{\phi}-\phi\|$, see Fig. 3 (a), which approaches zero with a pure GHZ state and error-free encoding of $\theta_{i}$. We similarly evaluate the distance for $\theta_{i}$, but with only outcomes of the $i$-node; this approach can recover $\theta_{i}$ using the non-private state. The distances for estimating $\phi$ are over an order of magnitude lower than for $\theta_{i}$, and remain constant with increasing measurement samples $\nu$, which shows $\phi$ was successfully estimated. We then evaluate the variance in the estimates for the $\phi$ as a function of $\nu$, see Fig. 3 (b), to obtain the precision scaling. The experimentally measured variances generally approach the scaling of $\nu^{-1}$, which is attainable with a pure GHZ state and optimal measurements, which confirms our estimate of $\phi$ is near optimal. We also plot the precision scaling for $\theta_{i}$ using the directly evaluated privacy parameter, $\varepsilon_{p}=0.005\pm 0.002$. This represents the upper bound on the precision scaling for $\theta_{i}$, based on the state quality, and is only realisable through an optimal (but generally unknown) measurement.
@@
Following the approach in Ref. [25] we will verify the privacy independently using the quantum Fisher information (QFI), a standard tool for benchmarking quantum metrology schemes, see Fig 4(c). The QFI quantifies the extractable information about a parameter encoded on a system with an optimal measurement. For a pure GHZ state and measuring in $X$ on all qubits, one can estimate $\phi$ with a theoretical QFI$=1$. Using the density matrices $\rho$ in our experiment we calculated the highest QFI=$0.90\pm 0.02$ when $\delta\tau=0$ while higher noise results in a lower value. We then evaluate the privacy parameter, $\varepsilon(\theta_{i})$, directly from $\rho$ for each local node value, $\theta_{i}$. We do this by calculating the reducible QFI of $\theta_{i}$, when encoded by the $i$-node and when the other $j\neq i$ nodes encode functions of $\theta_{i}$ such that the resulting QFI is minimised. For a GHZ state, the reducible QFI of any local parameter $\theta_{i}$ is zero, by letting the other nodes encode the local parameters $\theta_{j}=-\theta_{i}/(n-1)$, see Appendix 4 for details. For the non-private state $|+\rangle^{\otimes n}$, we obtain $\varepsilon_{p}=1/n^{2}$. We calculate $\varepsilon(\theta_{i})$ for all cases of $\delta\tau$, see Fig 4(c), we also evaluate the protocol privacy parameter, $\varepsilon_{p}\doteq\max_{i}\varepsilon(\theta_{i})$ for each case. The experimentally measured values for $\varepsilon_{p}$, of all tested $\delta\tau$, are smaller than the calculated value for the non-private state by over one order of magnitude—this confirms the privacy of the states we prepared.
@@ III Discussion
We demonstrated a quantum-private distributed sensing task involving three nodes that measure a global phase $\phi$ without revealing the local sensor values $\theta_{i}$—the precision in estimating $\phi$ is at least two orders of magnitude better than for any $\theta_{i}$. We assumed the availability of pairwise private channels between each node and the honest verifier for securely reporting outcomes in the verification protocol, this is a common requirement that is satisfied for privacy-enabled protocols [11, 10, 13, 15, 14]. Our work highlights a number of deficiencies in the implemented PPE protocol. First, we found the bounds on the fidelity of the GHZ state and privacy parameter of the task, which are obtained from the failure rate in the verification protocol, were exceedingly loose. Tighter security, for example by rethinking the privacy definition in a composable security framework, would make networked distributed sensing more resource efficient, and there is no fundamental reason preventing these from being derived in the future. Second, in its current form, the PPE protocol requires all copies of the shared states to be stored in memories before they can be used. The finite bandwidth even of multi-mode memories would strongly limit the number of copies that can be used, and as we have seen from the performance penalties of our finite round analysis, PPE would effectively become impractical. One may alleviate the need for large memories by introducing a trusted random variable that chooses whether a distributed state in a sequence is used or not for verification, such that each node only requires a single-qubit quantum memory [31]. Alternatively, one may consider GHZ-state certification protocols which exploit fast, low-loss optical switches to randomly select a copy out of an ensemble which can be used for parameter estimation, while the remaining copies are certified [32, 33, 34]. While the PPE protocol we demonstrated has some caveats, we also identified pathways for improvements, which, once addressed, will unlock a wider range and more efficient distributed sensing applications for quantum networks.
@@ Acknowledgements
We thank A. Pickston, L. Stroh, D. Markham and N. Shettell for helpful discussions. This work was supported by the UK Engineering and Physical Sciences Research Council (Grant Nos. EP/T001011/1.). FG acknowledges funding from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under the DFG Individual Grant BR2159/6-1.
@@ 4 Private parameter estimation protocols
2: For $i\in\{1,2,\dots,n\}$, do the following: (a) The Verifier selects $N_{t}/(2n)$ copies independently and uniformly at random. (b) For each selected copy, the Verifier instructs each party to perform the measurement corresponding to their part of the stabilizer operator $K_{i}$. (c) For each selected copy, the parties send their measurement outcome to the Verifier, who computes the outcome of the measurement of $K_{i}$. The copy passes the test if the measurement outcome of $K_{i}$ is +1. Let $N_{\text{pass},i}$ be the number of copies that pass the stabilizer test for $K_{i}$.

### v2

@@ I Introduction
Quantum networks allow connected nodes to perform tasks such as multi-user quantum cryptography [1, 2], distributed quantum computation [3, 4], and distributed quantum sensing [5, 6, 7, 8, 9]. The increased connectivity and sharing of resources over networks will likely prompt privacy-enabled protocols to protect sensitive information. A well known example of a private computing protocol is blind quantum computing, which allows clients to access remote quantum processors without revealing the computational algorithm to the servers [10, 3, 4]. Another established application is anonymous quantum communication, where users share cryptographic keys within a quantum network without revealing the identities of the keyholders to an eavesdropper, or other users within the network [11, 12, 13, 14, 15, 16]. Similar ideas from quantum cryptography can be incorporated into remote sensing tasks operating over quantum networks to provide both security and privacy features [17, 18, 19, 20].
Distributed quantum sensing typically involves spatially separated quantum sensors to monitor a global property, i.e., non-localised phenomena, such as magnetic fields, temperature sensing, or drifts between remote clocks [6]. The goal is to estimate a global value $\phi$ which is a linear function of local sensor values $\theta_{i}$ [21], such as the weighted sum—this differs from multi-parameter estimation where each $\theta_{i}$ is estimated, for which it is often optimal to use local strategies [22, 23]. In certain scenarios, distributed sensing can exploit entanglement generation in networks to estimate the global parameter directly with Heisenberg-limited precision scaling [17, 24, 8]. Additionally, entanglement can be used to guarantee security, ensuring only authorised users learn the estimated parameters, and privacy—delegating the sensing task to an untrusted node to perform measurements without learning the parameters [18, 19]. Recently, a two-user scheme has been proposed [25] and demonstrated [26] by sharing a Bell pair between two parties, one who performs the sensing task while only the other learns the estimated parameter.
@@
Here we consider multi-user quantum-private distributed sensing, a task known as private parameter estimation (PPE), for which security bounds have recently been established in Ref [27]. We illustrate the conceptual arrangement of the PPE task in Fig.1. The $n$ remote sensors are used to estimate a global phase, $\phi$, which is a sum $\phi\doteq\sum_{i}\theta_{i}$ of $\theta_{i}$ local phase values, with $i\in\{1,\dots,n\}$. The goal is to prevent information about the local $\theta_{i}$ from being learned by any node or eavesdroppers monitoring the communication. Analogous to quantum conference key agreement [28, 29, 16, 30], the correlations of the $n$-partite GHZ state, |$\ket{GHZ}\doteq(\ket{0}^{\otimes{n}}+\ket{1}^{\otimes{n}})/\sqrt{2}$ where |$\ket{0}$ and |$\ket{1}$ are computational basis states, are exploited to perform the distributed sensing task privately. The nodes must first verify that they are sharing the GHZ state to ensure they are performing the protocol in private and with integrity. Once this is achieved, the $\theta_{i}$ are encoded onto the shared GHZ state by coupling each node’s qubit to their respective quantum sensor, then each node measures their qubit and announces the result.
We now outline the state verification protocol [27]. One of the $n$ nodes is designated as the verifier, who is assumed to be honest and has pairwise private channels with each remaining node. An untrusted quantum server prepares and distributes $N_{t}$ copies of the $n$-partite GHZ state to all nodes who store them locally in quantum memories. The verifier randomly chooses half of the copies ($N_{t}/2$) and instructs the group to measure them according to one of the stabilizers, $K_{i}$. The set of stabilizers are given by cyclic permutations of $K_{i}=-X^{(1)}X^{(2)}\dots Y^{(i)}Y^{(i+1)}\dots X^{(n-1)}X^{(n)}$ for $i=1,2,\dots,n$ and $K_{i=n+1}=X^{\otimes n}$. By definition, the measurement outcomes of $K_{i}$ result in a $+1$ eigenvalue for the $n$-qubit GHZ state. Each node performs their respective measurement then communicates their outcome to the verifier directly using their private channels. The verifier computes a failure rate, $f$, as the fraction of copies corresponding to a $-1$ eigenvalue. The tests succeeds if $f\leq 1/(2n^{2})$ then one remaining untested copy is chosen at random as the target copy, $\rho_{j}$, for the parameter estimation task. This verification scheme is closely related to the one in Ref. [31], which is generalised to verify graph states using the full set of stabilizer measurements requiring $2n$ measurements, while in Ref. [27] only $n+1$ measurements are required for the GHZ state. This verification protocol guarantees a lower bound on the fidelity $F(\rho_{j},\ket{GHZ})$ of the target copy with the GHZ state up to some nominal probability,
@@
$\mathbb{P}(\textbf{A})$ is the probability that A is true, given the failure rate $f$, number of users $n$, and two positive variables $c$ and $m$ which relate to the statistical certainty based on the total number of copies, $N_{t}=\lceil 2mn^{5}\text{log}(n)\rceil$, used in the protocol (more details in the Appendix 2). Minimising $c$ improves the fidelity lower bound but decreases the probability of the inequality holding true. The parameter $m$ can be increased commensurately to offset the reduction by $c$, but this requires more test copies. Use of quantum memories in the verification step safeguards against an adversary swapping out non-test copies with non-private states, $|+\rangle^{\otimes{n}}$ where $|+\rangle\doteq(|0\rangle+|1\rangle)/\sqrt{2}$ which would allow the adversary to learn the local phases from the outcomes. We assume the future availability of such quantum memories and prepare multi-partite GHZ states in the lab to investigate the implementation of the verification scheme presented in Ref. [27].
@@ II Results
To implement the PPE protocol, the three nodes perform the verification protocol by measuring the stabilizers $K_{i}$, which we pick from a subset of the recorded QST measurements. For the optimal state we obtain an average failure rate $f=0.039\pm 0.005$ and from Ref [27] the verification step is successful if $f\leq 1/(2n^{2})$, in our three-node case this sets a threshold of $f\leq 0.05\dot{5}$ which our state satisfies. Using $f$ we establish a lower-bound fidelity of $0.769\pm 0.006$, via $F(\rho,GHZ)\geq 1-2nf$ when assuming infinite resource-state copies ($c\rightarrow 0$), which is smaller than the fidelity obtained via QST. For our three-node scenario it would be more optimal to use QST, however the measurements scale exponentially with $n$, while the verification protocol scales linearly requiring $n+1$ stabilizers. We also establish an upper-bounded privacy parameter, $\varepsilon_{p}\leq 1.3\pm 0.2$ using $\varepsilon_{p}\leq 24n^{-2}\sqrt{2nf}$, again assuming infinite copies, see Appendix 2 for details. For PPE, $\varepsilon_{p}=0$ corresponds to perfect privacy, wherein $\theta_{i}$ cannot be learned by any node nor an eavesdropper and $\phi$ can be estimated ideally [27]. Using the density matrix reconstructed from QST, we calculate the QFI (see below) and obtain $\varepsilon_{p}=0.005\pm 0.002$, which is much smaller than the upper bound from the verification protocol. We implement the phase sensing task for two global phases $\{\phi_{1},\phi_{2}\}$, see Fig. 3. To ensure each phase element encodes $\theta_{i}$ correctly we perform a full phase sweep for each, see Appendix 2 for details, then we independently set $\theta_{i}$ to prepare each $\phi$. To estimate $\phi$, all nodes measure in $X$ then compute the expectation value using all outcomes and evaluate $\phi=\textrm{cos}^{-1}[\langle X^{\otimes n}\rangle]$. The accuracy of the estimate $\hat{\phi}$ to the true value $\phi$ is evaluated as the distance, $\|\hat{\phi}-\phi\|$, see Fig. 3 (a), which approaches zero with a pure GHZ state and error-free encoding of $\theta_{i}$. We similarly evaluate the distance for $\theta_{i}$, but with only outcomes of the $i$-node; this approach can recover $\theta_{i}$ using the non-private state. The distances for estimating $\phi$ are over an order of magnitude lower than for $\theta_{i}$, and remain constant with increasing measurement samples $\nu$, which shows $\phi$ was successfully estimated. We then evaluate the variance in the estimates for the $\phi$ as a function of $\nu$, see Fig. 3 (b), to obtain the precision scaling. The experimentally measured variances generally approach the scaling of $\nu^{-1}$, which is attainable with a pure GHZ state and optimal measurements, which confirms our estimate of $\phi$ is near optimal. We also plot the precision scaling for $\theta_{i}$ using the directly evaluated privacy parameter, $\varepsilon_{p}=0.005\pm 0.002$. This represents the upper bound on the precision scaling for $\theta_{i}$, based on the state quality, and is only realisable through an optimal (but generally unknown) measurement.
@@
Following the approach in Ref. [27] we will verify the privacy independently using the quantum Fisher information, a standard tool for benchmarking quantum metrology schemes, see Fig 4(c). The QFI quantifies the extractable information about a parameter encoded on a system with an optimal measurement. For a pure GHZ state and measuring in $X$ on all qubits, one can estimate $\phi$ with a theoretical QFI${}=1$. Using the density matrices $\rho$ in our experiment we calculated the highest QFI ${}=0.90\pm 0.02$ when $\delta\tau=0$ while higher noise results in a lower value. We then evaluate the privacy parameter, $\varepsilon(\theta_{i})$, directly from $\rho$ for each local node value, $\theta_{i}$. We do this by calculating the reducible QFI of $\theta_{i}$, when encoded by the $i$-node and when the other $j\neq i$ nodes encode functions of $\theta_{i}$ such that the resulting QFI is minimised. For a GHZ state, the reducible QFI of any local parameter $\theta_{i}$ is zero, by letting the other nodes encode the local parameters $\theta_{j}=-\theta_{i}/(n-1)$, see Appendix 4 for details. For the non-private state $|+\rangle^{\otimes n}$, we obtain $\varepsilon_{p}=1/n^{2}$. We calculate $\varepsilon(\theta_{i})$ for all cases of $\delta\tau$, see Fig 4(c), we also evaluate the protocol privacy parameter, $\varepsilon_{p}\doteq\max_{i}\varepsilon(\theta_{i})$ for each case. The experimentally measured values for $\varepsilon_{p}$, of all tested $\delta\tau$, are smaller than the calculated value for the non-private state by over one order of magnitude—this confirms the privacy of the states we prepared.
@@ III Discussion
We have demonstrated a quantum-private distributed sensing task involving three nodes that measure a global phase $\phi$ without revealing the local sensor values $\theta_{i}$—the precision in estimating $\phi$ is at least two orders of magnitude better than for any $\theta_{i}$. We assumed the availability of pairwise private channels between each node and the honest verifier for securely reporting outcomes in the verification protocol, this is a common requirement that is satisfied for privacy-enabled protocols [12, 11, 14, 16, 15]. Our work highlights a number of outstanding challenges in the implemented PPE protocol. First, we found the bounds on the fidelity of the GHZ state and privacy parameter of the task, which are obtained from the failure rate in the verification protocol, were exceedingly loose. Tighter security, for example by rethinking the privacy definition in a composable security framework [33], would make networked distributed sensing more resource efficient, and there is no fundamental reason preventing these from being derived in the future. Second, in its current form, the PPE protocol requires all copies of the shared states to be stored in memories before they can be used. The finite bandwidth even of multi-mode memories would strongly limit the number of copies that can be used, and as we have seen from the performance penalties of our finite round analysis, PPE would effectively become impractical. One may alleviate the need for large memories by introducing a trusted random variable that chooses whether a distributed state in a sequence is used or not for verification, such that each node only requires a single-qubit quantum memory [34]. Alternatively, one may consider GHZ-state certification protocols which exploit fast, low-loss optical switches to randomly select a copy out of an ensemble which can be used for parameter estimation, while the remaining copies are certified [35, 36, 37]. Another improvement required to scale to more sensors is to revisit the current $1/n^{2}$ scaling for the fidelity bounds required for validation. Regarding the resource state, it was recently proven in Ref. [38] that the GHZ state, and equivalents up to a local unitary, is the only private state for estimating certain linear functions of inputs which our task falls within. While the PPE protocol we have demonstrated has some caveats, we also identified pathways for improvements, which, once addressed, will unlock a wider range and more efficient distributed sensing applications for quantum networks.
@@ Acknowledgements
We thank A. Pickston, L. Stroh, D. Markham and N. Shettell for helpful discussions. This work was supported by the UK Engineering and Physical Sciences Research Council (Grant Nos. EP/T001011/1.). FG acknowledges funding by the European Union’s Horizon Europe research and innovation program under the project “Quantum Security Networks Partnership” (QSNP, Grant Agreement No. 101114043) and by a French national quantum initiative managed by Agence Nationale de la Recherche in the framework of France 2030 with the reference ANR-22-PETQ-0009. FG did not contribute to this work on behalf of Leonardo S.p.A.
@@ 4 Private parameter estimation protocols
2: For $i\in\{1,2,\dots,(n+1)\}$, do the following: (a) The Verifier selects $N_{t}/(2n)$ copies independently and uniformly at random. (b) For each selected copy, the Verifier instructs each party to perform the measurement corresponding to their part of the stabilizer operator $K_{i}$. (c) For each selected copy, the parties send their measurement outcome to the Verifier, who computes the outcome of the measurement of $K_{i}$. The copy passes the test if the measurement outcome of $K_{i}$ is +1. Let $N_{\text{pass},i}$ be the number of copies that pass the stabilizer test for $K_{i}$.